OpenStudy (anonymous):

Can someone check my answer? x^2 + 5x + 6 = 0 (Points : 4) {2, 3} {-2, -3} {1,- 6} {-1, 6}<---- My answer

OpenStudy (anonymous):

see product is 6 nd sum is -5..........soo both shud be negative

OpenStudy (anonymous):

But thats not Choice

OpenStudy (anonymous):

a*

OpenStudy (lalaly):

(x+2)(x+3)=0 x=-2,-3

OpenStudy (anonymous):

x^2 + 5x + 6 = 0 x^2 + 2x + 3x + 6 = 0 x(x+2) + 3(x+2) (x+2)(x+3)

OpenStudy (anonymous):

So Is it he Second one?

OpenStudy (anonymous):

yea.........wen u got options.........do it this way sum is negative nd product is +ve so..........both shud be _ve.......u got nly 1 option there